OpenStudy (anonymous):

The Nut Barn requires 15 pounds of mixed walnuts and pecans from their wholesaler. Walnuts cost $2.26 per pound and pecans cost $1.96 per pound. How many pounds of pecans are needed if they only want to pay $2.18 per pound for the mixture?

hero (hero):

Hint: w + p = 15 2.26w + 1.96p = 2.18(15)

hero (hero):

Solve the system

OpenStudy (anonymous):

I have the answer I just want to make sure if its right peacans=4